# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v3

@@ A.1 RECAST-30K
To rigorously evaluate the ability of language models to follow complex, constraint-rich instructions, we construct a dataset consisting of 29,939 examples, which we refer to as RECAST-30K. RECAST-30K stands out for its high density and diversity of constraints, setting it apart from existing instruction datasets.
@@ A.1.1 Constraint Density
As discussed in Section 5.3, we also provide detailed statistics under different setups in Figures7, 8, and 9, showing distribution of datasets’ variants with different maximum constraint limits (5, 10, and 15).
@@ A.1.2 Constraint Types
To understand how constraint types are distributed in RECAST-30K, we visualize their distribution in Figure 10. The results suggest that RECAST-30K contains a highly diverse set of constraints that closely reflect real-world user instructions. Similarly, we provide results of the RECAST-30K variants with reduced sets of constraints. Figures 11, 12, and 13 visualize the constraint type distributions when the number of constraints is limited to 5, 10, and 15, respectively.
@@
A.1.3 Discussion on Realism and Diversity
@@
A.2 RECAST-Test
@@
To provide a comprehensive analysis of the constraint density distribution within the RECAST-Test, we conducted an in-depth examination of the constraint density across different subsets of RECAST-Test. Figures 14, 15, 16, and 17 present the constraint density distributions of the full RECAST-Test and its variants with a maximum constraints limit of 5, 10, 15, respectively. These results collectively demonstrate the richness of constraint in the RECAST-Test, which is crucial for evaluating the effectiveness of models’ ability to follow complex instructions. The detailed analysis of constraint density distribution across different subsets provides valuable insights into the dataset’s structure and utility.
@@
Similarly, Figures 18, 19, 20, 21 illustrates the constraint type distributions of the full RECAST-Test with all constraints included and its variants. These results demonstrate that RECAST-Test covers a highly diverse range of constraint types, effectively capturing the complexity and variability found in real-world instructions. The detailed breakdown of constraint type distributions across different variants offers valuable insights into the dataset’s structural diversity and practical utility.
@@
Appendix B Human Evaluation
@@
B.1 Evaluation for Constraint Filtering
@@
B.2 Evaluation for Optimal Instruction Selection
@@
B.3 Evaluation for Optimal Response Selection
@@
Appendix C Experimental Details
C.1 Baseline
@@
  • ConiferSun et al. [2024]: Constructs a GPT-4-refined dataset with hierarchical constraint instructions and employs curriculum tuning to enhance complex instruction-following.
@@
  • CrabQi et al. [2024]: Enhances LLMs by back-translating model responses into constraint-rich instructions using a stronger teacher model.
@@
  • I-SHEEPLiang et al. [2024b]: Aligns LLMs from scratch via iterative self-generated instruction–response tuning without external supervision.
@@
  • MUFFINLou et al. [2023]: Diversifies instruction data by scaling input facets per task, yielding more robust multi-faceted alignment.
@@
  • ShareGPT555huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered: Provides real user–assistant dialogues as a large-scale resource for training conversational instruction-following.
@@
  • Evol-InstructXu et al. [2024]: Generates complex instructions through iterative rewriting, producing progressively harder instruction–response pairs.
@@
  • SuriPham et al. [2024]: Back-translates long-form texts into multi-constraint prompts, improving structured generation over extended contexts.
@@
  • Tülu 3 Persona IFLambert et al. [2024]: Synthesizes constraint-following data from diverse persona prompts to enhance controllable instruction tuning.
C.2 General Capability Evaluation Benchmark
@@
We employ six diverse benchmarks spanning different aspects of language understanding and reasoning:
@@
  • FollowBench Jiang et al. [2024b] comprehensively evaluates instruction-following capability through five types of fine-grained constraints (Content, Situation, Style, Format, and Example) using a multi-level mechanism that incrementally adds constraints to initial instructions. This benchmark systematically measures models’ ability to satisfy multiple constraints simultaneously, with difficulty levels ranging from simple single-constraint instructions to complex multi-constraint scenarios. The evaluation employs metrics including Hard Satisfaction Rate (HSR) for complete constraint satisfaction and Soft Satisfaction Rate (SSR) for partial compliance.
@@
  • IFEval Zhou et al. [2023] assesses models’ ability to follow explicit formatting instructions through strict, verifiable metrics rather than subjective content evaluation. The benchmark focuses on precise adherence to specific requirements such as keyword inclusion, format specifications, and structural constraints. This allows for rigorous, code-based evaluation that eliminates ambiguity in measuring instruction compliance.
  • BBH Suzgun et al. [2023] comprises 23 carefully selected challenging tasks from the BigBench dataset that test diverse cognitive capabilities including multistep arithmetic, algorithmic reasoning, language understanding, and world knowledge. These tasks are chosen for their high difficulty, objective metrics, and strong correlation with human preferences. The benchmark provides insights into models’ reasoning abilities across domains ranging from boolean expression evaluation to sarcasm detection.
@@
  • GPQA Rein et al. [2023] features expert-crafted questions from PhD-level domain specialists in biology, physics, and chemistry designed to be challenging for non-experts yet accessible to specialists. The dataset undergoes rigorous validation to ensure both factual accuracy and appropriate difficulty levels. Access restrictions and gating mechanisms protect against data contamination, maintaining the benchmark’s integrity for evaluating advanced scientific knowledge.
@@
  • MuSR Sprague et al. [2024] tests models’ ability to integrate reasoning with long-context understanding through algorithmically generated complex problems averaging 1,000 words in length. The benchmark includes murder mysteries, object placement puzzles, and team allocation optimization tasks that require maintaining coherent reasoning across extended contexts. Few models achieve above-random performance, making it an effective discriminator of advanced reasoning capabilities.
  • MMLU-PRO Wang et al. [2024b] represents an enhanced version of the original MMLU benchmark, addressing previous limitations through increased difficulty and improved quality control. The benchmark presents 10 answer choices instead of 4, requires deeper reasoning, and underwent expert review to eliminate noisy or unanswerable questions. This refinement results in a more challenging and reliable assessment of multitask language understanding across diverse knowledge domains.
C.3 Details of Data Generation
@@ C.4 Details of SFT
This section details our model training configuration based on the LlamaFactory Zheng et al. [2024] framework. We employed Supervised Fine-Tuning (SFT) with a maximum sequence length of 4096 tokens. For optimization, we used a linear learning rate scheduler with a peak learning rate of 2.0e-5, 3% warmup ratio, and trained for 3 epochs. We configured training with a per-device batch size of 4 and gradient accumulation steps of 32, resulting in an effective batch size of 128. To optimize training efficiency, we utilized DeepSpeed with ZeRO stage 3 optimization and bfloat16 mixed precision. All experiments were conducted on a cluster of 8 NVIDIA H800 GPUs.
C.5 Details of RLVC
@@
For reinforcement learning, we implemented our RLVC approach based on the VeRL Sheng et al. [2024] framework with Group Relative Policy Optimization (GRPO). Our configuration used a learning rate of 1e-6 with maximum sequence lengths of 1024 tokens for both prompts and responses. We employed a batch size of 512 with PPO mini-batches of 128 and micro-batches of 16 per GPU. For optimization stability, we incorporated KL divergence regularization with a coefficient of 0.001 using the low-variance KL implementation, while enabling gradient checkpointing for memory efficiency. The rollout process utilized tensor model parallelism with a size of 2 and vLLM for acceleration, generating 16 candidate responses per prompt with 60% GPU memory utilization. We integrated our custom constraint-specific reward functions through the thread-based reward manager with asynchronous reward calculation. All experiments were conducted on 8 GPUs for 200 optimization steps.
Appendix D Constraint Taxonomy
@@
Prior to generating constraints, we constructed a comprehensive constraint taxonomy composed of two categories: rule-based constraints and model-based constraints. The constraint categories include 20 types, with definitions and representative examples shown in Table 6 and Table 7. These types serve as the basis for generating diverse and realistic constraints.
D.1 Rule-based Constraints
@@
Rule-based constraints are objective requirements that can be verified using deterministic methods, such as Python functions. These constraints are structural, lexical, or quantitative in nature, such as paragraph count, keyword inclusion, and word limits. To generate these constraints, we implemented eight rule-based extractors that identify verifiable properties from text. These extractors analyze syntactic patterns, keyword frequency, numerical values, and other measurable elements to produce concrete constraint instances.
@@
D.2 Model-based Constraints
Model-based constraints are subjective and require semantic interpretation or qualitative judgment. These involve higher-level aspects such as style, tone, emotion, and content relevance. We developed a taxonomy of 15 constraint types encompassing semantic fidelity, stylistic nuance, tonal variation, and domain-specific content requirements. These were used to guide the LLM in generating rich and context-sensitive constraint instances for each response.
@@ D.2 Model-based Constraints
D.3 Templates for Rule-based Constraint Generation
@@
D.3.1 Templates for Length Constraints (Word Level)
@@
D.3.2 Templates for Length Constraints (Sentence Level)
@@
D.3.3 Templates for Other Constraints
@@
Appendix E Main Prompts
@@
Appendix F Limitations
Despite the promising results of our work, several limitations warrant acknowledgment. First, the computational cost of constraint verification presents a significant challenge, particularly for model-based constraints that require LLM-based evaluation during reinforcement learning. This verification overhead may limit the scalability of our approach in resource-constrained environments. Second, our approach assumes constraints can be verified independently, which may not always hold when constraints interact with or contradict each other—a common occurrence in complex real-world scenarios. Third, the effectiveness of RLVC depends on the quality of constraint verification mechanisms, which themselves have inherent limitations, particularly for subjective constraints where even human evaluators might disagree. Finally, our evaluation primarily focuses on English, and the approach may require adaptation for other languages with different linguistic structures and constraint manifestations. Future work could address these limitations through more efficient verification methods, improved handling of constraint interactions, and cross-lingual evaluations.

### v4

@@ B.2 RECAST-30K
To rigorously evaluate the ability of language models to follow complex, constraint-rich instructions, we construct a dataset consisting of 29,939 examples, which we refer to as RECAST-30K. RECAST-30K stands out for its high density of constraints, setting it apart from existing instruction datasets.
@@ B.2.1 Constraint Density
As discussed in Appendix D.6, we also provide detailed statistics under different setups in Figures5, 6, and 7, showing distribution of datasets’ variants with different maximum constraint limits (5, 10, and 15).
@@ B.2.2 Constraint Types
To understand how constraint types are distributed in RECAST-30K, we visualize their distribution in Figure 8. The results suggest that RECAST-30K contains a highly diverse set of constraints that closely reflect real-world user instructions. Similarly, we provide results of the RECAST-30K variants with reduced sets of constraints. Figures 9, 10, and 11 visualize the constraint type distributions when the number of constraints is limited to 5, 10, and 15, respectively.
@@
B.2.3 Discussion on Realism and Diversity
@@
B.3 Cost Analysis of Constructing RECAST-30K
To provide transparency on the resource requirements of building our dataset, we report the estimated costs incurred at each stage of the RECAST pipeline. The total expenditure was approximately \$175, broken down as shown in Table 6.
Table 6: Estimated cost of constructing RECAST-30K across different pipeline stages.
Pipeline Stage Estimated Cost (USD)
Rule-based Constraints Construction \$0 (no API calls required)
Model-based Constraints Construction \$70
Instruction Enhancement \$85
Majority Voting \$20
Total \$175
These figures demonstrate that RECAST-30K can be constructed at relatively low cost, underscoring the scalability and practicality of our pipeline for broader adoption in future instruction-following research.
B.4 RECAST-Test
@@
To provide a comprehensive analysis of the constraint density distribution within the RECAST-Test, we conducted an in-depth examination of the constraint density across different subsets of RECAST-Test. Figures 12, 13, 14, and 15 present the constraint density distributions of the full RECAST-Test and its variants with a maximum constraints limit of 5, 10, 15, respectively. These results collectively demonstrate the richness of constraint in the RECAST-Test, which is crucial for evaluating the effectiveness of models’ ability to follow complex instructions. The detailed analysis of constraint density distribution across different subsets provides valuable insights into the dataset’s structure and utility.
@@
Similarly, Figures 16, 17, 18, 19 illustrates the constraint type distributions of the full RECAST-Test with all constraints included and its variants. These results demonstrate that RECAST-Test covers a highly diverse range of constraint types, effectively capturing the complexity and variability found in real-world instructions. The detailed breakdown of constraint type distributions across different variants offers valuable insights into the dataset’s structural diversity and practical utility.
@@
Appendix C Human Evaluation
@@
C.1 Evaluation for Constraint Filtering
@@
C.2 Evaluation for Optimal Instruction Selection
@@
C.3 Evaluation for Optimal Response Selection
@@
Appendix D Experimental Details
D.1 Baseline
@@
  • Conifer (Sun et al., 2024): Constructs a GPT-4-refined dataset with hierarchical constraint instructions and employs curriculum tuning to enhance complex instruction-following.
@@
  • Crab (Qi et al., 2024): Enhances LLMs by back-translating model responses into constraint-rich instructions using a stronger teacher model.
@@
  • I-SHEEP (Liang et al., 2024): Aligns LLMs from scratch via iterative self-generated instruction–response tuning without external supervision.
@@
  • MUFFIN (Lou et al., 2023): Diversifies instruction data by scaling input facets per task, yielding more robust multi-faceted alignment.
@@
  • ShareGPT [^5] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered: Provides real user–assistant dialogues as a large-scale resource for training conversational instruction-following.
@@
  • Evol-Instruct (Xu et al., 2024): Generates complex instructions through iterative rewriting, producing progressively harder instruction–response pairs.
@@
  • Suri (Pham et al., 2024): Back-translates long-form texts into multi-constraint prompts, improving structured generation over extended contexts.
@@
  • Tülu 3 Persona IF (Lambert et al., 2024): Synthesizes constraint-following data from diverse persona prompts to enhance controllable instruction tuning.
D.2 General Capability Evaluation Benchmark
@@
We employ four diverse benchmarks spanning different aspects of language understanding and reasoning:
@@
  • FollowBench (Jiang et al., 2024b) comprehensively evaluates instruction-following capability through five types of fine-grained constraints (Content, Situation, Style, Format, and Example) using a multi-level mechanism that incrementally adds constraints to initial instructions. This benchmark systematically measures models’ ability to satisfy multiple constraints simultaneously, with difficulty levels ranging from simple single-constraint instructions to complex multi-constraint scenarios. The evaluation employs metrics including Hard Satisfaction Rate (HSR) for complete constraint satisfaction and Soft Satisfaction Rate (SSR) for partial compliance.
@@
  • IFEval (Zhou et al., 2023) assesses models’ ability to follow explicit formatting instructions through strict, verifiable metrics rather than subjective content evaluation. The benchmark focuses on precise adherence to specific requirements such as keyword inclusion, format specifications, and structural constraints. This allows for rigorous, code-based evaluation that eliminates ambiguity in measuring instruction compliance.
@@
  • GPQA (Rein et al., 2023) features expert-crafted questions from PhD-level domain specialists in biology, physics, and chemistry designed to be challenging for non-experts yet accessible to specialists. The dataset undergoes rigorous validation to ensure both factual accuracy and appropriate difficulty levels. Access restrictions and gating mechanisms protect against data contamination, maintaining the benchmark’s integrity for evaluating advanced scientific knowledge.
@@
  • MuSR (Sprague et al., 2024) tests models’ ability to integrate reasoning with long-context understanding through algorithmically generated complex problems averaging 1,000 words in length. The benchmark includes murder mysteries, object placement puzzles, and team allocation optimization tasks that require maintaining coherent reasoning across extended contexts. Few models achieve above-random performance, making it an effective discriminator of advanced reasoning capabilities.
D.3 Details of Data Generation
@@ D.4 Details of SFT
This section details our model training configuration based on the LlamaFactory (Zheng et al., 2024) framework. We employed Supervised Fine-Tuning (SFT) with a maximum sequence length of 4096 tokens. For optimization, we used a linear learning rate scheduler with a peak learning rate of 2.0e-5, 3% warmup ratio, and trained for 3 epochs. We configured training with a per-device batch size of 4 and gradient accumulation steps of 32, resulting in an effective batch size of 128. To optimize training efficiency, we utilized DeepSpeed with ZeRO stage 3 optimization and bfloat16 mixed precision. All experiments were conducted on a cluster of 8 NVIDIA H800 GPUs.
D.5 Details of RLVC
Policy Optimization
We employ Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our policy learning algorithm. GRPO generates multiple candidate responses for the same instruction and computes advantage estimates through intra-group comparisons, allowing the model to learn the relative quality differences among various responses to the same instruction.
During policy updates, GRPO maximises the following objective:
where:
Implementation Details
@@ D.5 Details of RLVC
For reinforcement learning, we implemented our RLVC approach based on the VeRL (Sheng et al., 2025) framework with Group Relative Policy Optimization (GRPO). Our configuration used a learning rate of 1e-6 with maximum sequence lengths of 1024 tokens for both prompts and responses. We employed a batch size of 512 with PPO mini-batches of 128 and micro-batches of 16 per GPU. For optimization stability, we incorporated KL divergence regularization with a coefficient of 0.001 using the low-variance KL implementation, while enabling gradient checkpointing for memory efficiency. The rollout process utilized tensor model parallelism with a size of 2 and vLLM for acceleration, generating 16 candidate responses per prompt with 60% GPU memory utilization. We integrated our custom constraint-specific reward functions through the thread-based reward manager with asynchronous reward calculation. All experiments were conducted on 8 GPUs for 200 optimization steps.
Training Dynamics
To better understand the optimization behaviour of RLVC, we track the evolution of both the model-based and rule-based rewards, together with their corresponding satisfaction rates, during the early phase of RL training. Table 8 reports the recorded values at steps 0, 50, 100, 150 and 200.
Table 8: Reward metrics and constraint satisfaction rates during the first 200 RLVC optimization steps.
Step 0 50 100 150 200
Model-based reward 0.411 0.428 0.445 0.431 0.415
MSR (%) 54.0 56.5 57.5 56.0 55.5
Rule-based reward 0.011 0.022 0.016 0.030 0.018
RSR (%) 4.8 5.5 5.4 5.8 10.9
The observed dynamics can be characterised in three phases. First, an early surge (0–100 steps): the model rapidly increases its model-based reward as it captures a set of relatively easy model-based constraints, accompanied by gradual improvements in rule-based satisfaction. Second, a slow-down phase (after $\sim$100 steps): once these low-hanging model-based objectives are largely addressed, optimization attention shifts toward more difficult, often rule-based objectives, causing the model-based reward to plateau or decline slightly, while rule-based metrics begin to improve more noticeably. Third, the overall trend: despite the plateau in the model-based reward, both MSR and RSR show mild upward trajectories, indicating that the model is improving its aggregated ability to satisfy an increasingly challenging mixture of constraints.
Notably, the scale of improvements differs across metric types: model-based reward increases modestly by about 88% relative (from $0.411$ to $0.445$ at peak), whereas rule-based reward—though starting at a much lower absolute value—doubles in magnitude and its satisfaction rate (RSR) grows from 4.84.8% to 10.910.9%. This contrast highlights the higher difficulty but also the meaningful progress in handling rule-based objectives. These dynamics suggest that RLVC first exploits readily learnable, model-evaluable constraints before reallocating capacity to harder objectives; consequently, monitoring both per-constraint-type rewards and aggregated metrics is important for diagnosing progress and for informing choices such as reward balancing, curriculum scheduling, or extended training.
D.6 Ablations and Analysis
Impact of constraint type during complex instruction learning.
Impact of constraint number during complex instruction learning.
Impact of specific components of RECAST.
To isolate component contributions, we evaluated RECAST variants with specific elements removed(Table 9). Complete RECAST-30k consistently outperforms all ablated configurations. Instruction-Only Enhancement exhibits the most significant performance degradation (12.2112.21% and 17.3717.37% decreases for Llama and Qwen), revealing that constraint-augmented responses without corresponding response modifications create instruction-response inconsistencies that confuse the model, substantially hampering performance. Response-Only Enhancement performs comparably to Without RECAST, suggesting that constraint specification in instructions drives most performance gains. These results confirm that RECAST’s complete instruction-response synthesis pipeline is essential for optimal constraint satisfaction capabilities.
  Configuration Average
Llama-3.1-8B Qwen2.5-7B
  RECAST-30K $30.54$ $31.25$
Response-Only Enhancement $27.00$ $25.54$
Instruction-Only Enhancement $18.33$ $13.88$
Without RECAST $26.96$ $26.83$
Appendix E Constraint Taxonomy
@@
Prior to generating constraints, we constructed a comprehensive constraint taxonomy composed of two categories: rule-based constraints and model-based constraints. The constraint categories include 19 types, with definitions and representative examples shown in Table 10 and Table 11. These types serve as the basis for generating diverse and realistic constraints.
E.1 Rule-based Constraints
@@
Rule-based constraints are objective requirements that can be verified using deterministic methods, such as Python functions. These constraints are structural, lexical, or quantitative in nature, such as paragraph count, keyword inclusion, and word limits. To generate these constraints, we implemented nine rule-based extractors that identify verifiable properties from text. These extractors analyze syntactic patterns, keyword frequency, numerical values, and other measurable elements to produce concrete constraint instances.
@@
E.2 Model-based Constraints
Model-based constraints are subjective and require semantic interpretation or qualitative judgment. These involve higher-level aspects such as style, tone, emotion, and content relevance. We developed a taxonomy of 10 constraint types encompassing semantic fidelity, stylistic nuance, tonal variation, and domain-specific content requirements. These were used to guide the LLM in generating rich and context-sensitive constraint instances for each response.
@@ E.2 Model-based Constraints
E.3 Templates for Rule-based Constraint Generation
@@
E.3.1 Templates for Length Constraints (Word Level)
@@
E.3.2 Templates for Length Constraints (Sentence Level)
@@
E.3.3 Templates for Other Constraints
@@
Appendix F Main Prompts
@@
LLM Usage Statement
LLMs did not contribute to the research ideation, experimental design, analysis, or manuscript writing. All conceptual and textual contributions are solely attributable to the authors.